NEED HELP ASAP PLEASE! (I already answered 1-4)

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 62 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $ 3,614,400 $ 3,055,100
Net income 817,600 625,700
Total $4,432,000 $ 3,680,800
Dividends:
On preferred stock $ 10,500 $ 10,500
On common stock 55,900 55,900
Total dividends $ 66,400 $ 66,400
Retained earnings, December 31 $ 4,365,600 $ 3,614,400
Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $ 5,101,970 $ 4,700,720
Cost of goods sold 1,696,520 1,560,800
Gross profit $ 3,405,450 $ 3,139,920
Selling expenses $ 1,194,050 $ 1,450,030
Administrative expenses 1,017,160 851,600
Total operating expenses $2,211,210 $2,301,630
Income from operations $ 1,194,240 $ 838,290
Other revenue 62,860 53,510
$ 1,257,100 $ 891,800
Other expense (interest) 328,000 180,800
Income before income tax $ 929,100 $ 711,000
Income tax expense 111,500 85,300
Net income $ 817,600 $ 625,700
Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   20Y2    20Y1
Assets
Current assets
Cash $ 924,380 $ 797,480
Marketable securities 1,399,060 1,321,530
Accounts receivable (net) 905,200 854,100
Inventories 686,200 525,600
Prepaid expenses 174,888 159,500
Total current assets $ 4,089,728 $ 3,658,210
Long-term investments 1,933,912 467,256
Property, plant, and equipment (net) 5,330,000 4,797,000
Total assets $ 11,353,640 $ 8,922,466
Liabilities
Current liabilities $ 1,278,040 $ 1,438,066
Long-term liabilities:
Mortgage note payable, 8% $ 1,840,000 $ 0
Bonds payable, 8% 2,260,000 2,260,000
Total long-term liabilities $ 4,100,000 $ 2,260,000
Total liabilities $ 5,378,040 $ 3,698,066
Stockholders' Equity
Preferred $0.70 stock, $50 par $ 750,000 $ 750,000
Common stock, $10 par 860,000 860,000
Retained earnings 4,365,600 3,614,400
Total stockholders' equity $ 5,975,600 $ 5,224,400
Total liabilities and stockholders' equity $ 11,353,640 $ 8,922,466

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

1. Working capital $ 2,811,688
2. Current ratio 3.2
3. Quick ratio 2.5
4. Accounts receivable turnover 11.6
5. Number of days' sales in receivables days
6. Inventory turnover
7. Number of days' sales in inventory days
8. Ratio of fixed assets to long-term liabilities
9. Ratio of liabilities to stockholders' equity
10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield
Answer #1

4. Accounts receivable turnover = Sales / Avg.Receivables

                                              = 5,101,970 / ( 905,200 +854100)/2 = 5101970 / 879650 = 5.8

5. Number of days' sales in receivables = 365 / Receivables Turnover = 365 / 5.8 = 62.93 Days

6. Inventory turnover = Cost of goods sold / Avg.Inventory

                              = 1,696,520 / ( 686,200 + 525600 ) / 2 = 1696520 / 605900 = 2.8

7. Number of days' sales in inventory = 365 / Inventory turnover = 365 / 2.8 = 130.36 Days

8. Ratio of fixed assets to long-term liabilities = 5,330,000 /4100000 = 1.3

9. Ratio of liabilities to stockholders' equity = 5,378,040 /5975600 = 0.9

10. Times interest earned = Income from operations / Interest Expense = 1,194,240 /328000 = 3.64

11. Asset turnover = Sales / Avg. Total Assets = 5,101,970 / ( 11,353,640 + 8922466 ) /2 = 5101970 / 10138053 = 0.503

12. Return on total assets = Net Income / Avg. Total Assets = 817,600 / 10138053 = 8.06%

13. Return on stockholders’ equity = Net Income / Avg. Total Assets

                                                  = 817,600 / ( 5,975,600 + 5224400)/2 = 817600 / 5600000 = 14.6%

14. Return on common stockholders’ equity = (Net Income - preferred dividends)/ Avg. common stockholders equity)

                                                  = ( 817600 - 10500 ) / ( 860000 + 860000 + 4,365,600 + 3,614,400 ) / 2

                                                  = 807100 / 4850000 = 16.64%

15. Earnings per share on common stock = (Net Income - preferred dividends) / Outstanding Common Stock

                                                   = 807100 / 86000 = $9.38

16. Price-earnings ratio = Market Price / EPS = 62 / 9.38 = 6.61

17. Dividends per share of common stock = Dividend on common stock / Outstanding Common Stock

                                                             = 55900 / 86000 = $0.65

18. Dividend yield = Dividend per share / Market Price = $0.65 / $62 = 1.048%